OpenStudy (anonymous):

Solve the system using the method of your choice: y = 3x – 4 2x + 4y = 26

OpenStudy (texaschic101):

substitution ok with you ?

OpenStudy (texaschic101):

y = 3x - 4.....so we sub in 3x - 4 for y back into the other equation 2x + 4y = 26 2x + 4(3x - 4) = 26 --- distribute through the parenthesis 2x + 12x - 16 = 26 --- add 16 from both sides 2x + 12x = 26 + 16 -- combine terms 14x = 42 x = 42/14 x = 3 now sub 3 in for x in either equation to find y y = 3x - 4 y = 3(3) - 4 y = 9 - 4 y = 5 we can even check our answers 2x + 4y = 26 2(3) + 4(5) = 26 6 + 20 = 26 26 = 26 (correct) so x = 3 and y = 5....or (3,5)

OpenStudy (anonymous):

thank you!:)